# Exploring a New Design Paradigm for Omnidirectional MAVs for Minimal Actuation and Internal Force Elimination: Theoretical Framework and Control

### v1

@@ III Study of Attainable Equilibria and
In this section we obtain the set of forced equilibria of the dynamics (4) and investigate their properties in terms of the corresponding four thrust forces produced at these configurations. We begin by giving some preliminary definitions. Afterwards, the main result is summarized in Proposition 1 whose proof is found in Sec. VII-B.
The value of a variable at an equilibrium is identified by a subscript $e$. Imposing $\dot{x}{(t)}=0$ in (4) implies $({x}_{2}(t),\dot{q}_{5}(t))=0$, which defines an equilibrium condition at which the state and the input become, respectively, $x_{e}=(x_{1e},0,{q}_{5e})^{T}$ and $u_{e}=(f_{e}\,\,u_{ve})^{T}$ with $f_{e}=(f_{2e}\,\,f_{3e}\,\,f_{4e}\,\,f_{5e})^{T}$ and $u_{ve}=(u_{v_{1e}}\,\,u_{v_{2e}})^{T}$. This condition yields $h(x_{1}(t),{x}_{2}(t))=0$ since it depends quadratically on the configuration velocity state ${x}_{2}(t)$, as apparent from Table I. Similarly, the viscous friction $\bar{\tau}_{f}(x_{22}(t))=0$. For PJD1, this reduces (4) to this static balance vector equation whose admissible solutions are the equilibrium state and input pairs $(x_{e},u_{e})$ such that
$g_{b}(C^{0}_{1e},q_{te})=W^{0}_{app,b}(f_{e},q_{te},q_{5e}),$ (6)
$g_{t}(C^{0}_{1e},q_{te})=W^{0}_{app,t}(q_{te},f_{2e},f_{3e},f_{4e}):=(\,\,D_{t}\,\,0_{6\times 3}\,\,)u_{e}$
$\quad\quad\quad\quad\quad=(\,\,\tau_{d2}\,\,0\,\,\tau_{d3}\,\,0\,\,\tau_{d4}\,\,0\,\,)_{6\times 1}^{T},\quad 0_{2}=u_{ve}$ (7)
where $\tau_{dj}=k_{j}f_{je}\sin(q_{j2e})$ – with $j=2,3,4$ – represents the drag component of the propeller of $B_{j}$ factorized on the axis $e_{j1}$ of the joint variable $q_{j1}$ at an equilibrium; $C^{0}_{1e}$ and $q_{te}=\operatorname{col}(\,q_{je})$, with $q_{je}=\operatorname{col}(\,q_{jie})$ $\forall j=2,3,4\text{ and }\forall i=1,2$, symbolizing an equilibrium of the base configuration $C^{0}_{1}$ and the passive joints configuration $q_{t}$, respectively.
@@ III-A Effect of the Aerodynamic Drag Torque at the Equilibrium
The presence of the torque $\tau_{dj}$ in (7) leads to the important consequence that the value of the joints $q_{te}$ at the equilibrium is not solely determined by the gravitational torque at equilibrium $g_{te}$, as it is the case for the 2D MAV presented in  [6], but also by a component of the equilibrium thrust $f_{e}$ at equilibrium.
This residual torque can be eliminated at any equilibrium by the spring introduced in PJD2. In fact, we can tune each torsion spring to have a variable stiffness $k_{s}:=\frac{\partial\tau_{sj}}{\partial q_{j2}}=k_{j}f_{je}\cos(q_{j2})$ where $\tau_{sj}$ is the torque absorbed by the spring at the joint angle $q_{j2}$. This task is facilitated by the fact that the term $k_{j}f_{je}$ is always constant since $f_{je}$ is found to be constant for any equilibrium of PJD2, as proven in Sec. VII-B (49). As a result, this component $\tau_{dj}$ is canceled at any PJD2 equilibrium, i.e., $\tau_{dj}=-\tau_{sje}$, transforming (7) to
$g_{t}(C^{0}_{1e},q_{te})=0_{6},\,\,\,u_{ve}=0_{2}.$ (8)
We solve (7) and (8) in Sec. VII-B for the joint equilibrium $q_{je}$. To distinguish between a joint equilibrium of PJD1 and PJD2, let $q_{te}^{[i]}$ denote a joint equilibrium solution of PJDi, with $i=1,2$, and $f^{[i]}_{e}$ be the associated thrust, respectively.
@@ III-B Minimal Thrust at Equilibrium vs Presence of Internal Forces
Before presenting the result on the equilibria property of the proposed MAV (4) we provide two definition for the thrust magnitudes at the equilibrium, in the case in which they are minimal and in the case in which there is internal force.
@@
$\mathcal{F}_{\rm min}^{\rm eq}:=\bigg\{f=col(f_{i})\in\mathbb{R}_{\geq 0}^{l}\bigg|\,\sum^{l}_{i=1}{f}_{i}=m_{\rm tot}\,g\bigg\}$ (9)
where $m_{\rm tot}$ is the total mass of the system. Instead, $f$ is said to have internal forces at the equilibrium if it belongs to $\mathcal{F}_{\rm int.for.}^{\rm eq}$, defined as:
$\mathcal{F}_{\rm int.for.}^{\rm eq}:=\bigg\{f=col(f_{i})\in\mathbb{R}_{\geq 0}^{l}\bigg|\,\sum^{l}_{i=1}{f}_{i}>m_{\rm tot}\,g\bigg\}$ (10)
@@
In other words, a minimal thrust magnitudes vector $f\in\mathcal{F}_{\rm min}^{\rm eq}$ has the least amount of total thrust magnitudes necessary to counteract the gravity effect at the equilibrium, whereas $f\in\mathcal{F}_{\rm int.for.}^{\rm eq}$ means that the vehicle generates more total thrust than what is strictly necessary for compensating gravity, thus having some wasted internal force.
Moreover, there exists a unique orientation for each propeller in the system such that that when all the propellers assume such orientation the absence of internal force is guaranteed at hovering. The following lemma provides the formal statement.
Lemma 1.
Consider a vehicle with a thrust magnitudes vector $f=col(f_{i})$ $\in\mathbb{R}_{\geq 0}^{l}$ with the corresponding propellers configurations $C^{0}_{i}(C^{0}_{1},\alpha_{i})\in SE(3)$, relative to a right-handed inertial frame whose z-axis direction opposes the axis along which the gravitational force is applied, where $C^{0}_{1}\in SE(3)$ is the base configuration and $\alpha_{i}\in\mathbb{S}^{2}$ is the unit direction vector of the propeller $i$ expressed in the inertial frame. At any equilibrium for the MAV with a given $\bar{C}^{0}_{1}$, there is no internal force generated at such equilibrium if and only if all propellers, whose thrust magnitudes are $f_{i}>0$, are pointing in the direction of inertial frame’s z-axis, i.e. $C^{0}_{i}(\bar{C}^{0}_{1},\alpha_{i})\,e_{3}=e_{3}$ $\forall i\in\{1,\dots,l\}$ with $e_{3}=(\,0\,0\,1\,0\,)^{T}$.
Proof.
See Sec. VII-A. ∎
The following proposition establishes the main result of this section.
@@
    For any base pose $C^{0}_{1}\in SE(3)$, there exists an equilibrium for PJD2 in which the joint equilibrium configuration $q_{te}^{[2]}$ satisfies $C^{0}_{j}(C^{0}_{1e},q^{[2]}_{je})e_{3}=e_{3}$ $\forall j\in\{2,\dots,5\}$ and a thrust magnitudes vector $f^{[2]}_{e}\in\mathcal{F}_{\rm min}^{\rm eq}$ is produced whose elements are constant $\forall C^{0}_{1}$ and equal to (49).
@@
    For any base pose $C^{0}_{1}$, PJD2 has no internal force at any equilibrium $q_{te}^{[2]}$, by Definition 1, as a corollary of $f^{[2]}_{e}\in\mathcal{F}_{\rm min}^{\rm eq}$ (Statement 1).
@@
    When $\tau_{dj}\neq 0$, PJD1 has no joint equilibrium in which each propeller produces a minimal thrust: $C^{0}_{j}(C^{0}_{1e},q^{[1]}_{je})e_{3}\neq e_{3},\,\forall\tau_{dj}\neq 0$.
@@
    When $\tau_{dj}\neq 0$, it follows from Statement 3 and Lemma 1 that PJD1 has an internal force $f^{[1]}_{e}\in\mathcal{F}_{\rm int.for.}^{\rm eq}$. The amount of internal force changes with the parameters $k_{i}$,c\,c and $m_{p}$, defined in Table II.
@@
See Sec. VII-B. ∎
The state equilibria sets for PJD1 and PJD2 can thus be defined as $\mathbb{D}^{[1]}_{x_{e}}$ and $\mathbb{D}^{[2]}_{x_{e}}$, respectively, with $i=1,2$
[…]
while the associated thrust equilibrium sets are given by
[…]
We conclude the section remarking that the perfect vertical alignment of propellers in any PJD2 equilibrium is possible thanks to each spring torque $\tau_{sje}$ absorbing the residual drag $\tau_{dj}$ affecting the joint variable $q_{j1}$. This allows any joint equilibrium $q_{te}^{[2]}$ to be dictated only by the gravity acting on the CoMs of the pendulum-like links B2-B4, given that the pivots, i.e., the joints, are passive.
@@
The design using PJD1’s constitutes still an excellent compromise between mechanical simplicity and energy efficiency, with the ability to outperform the state of the art designs when it comes to the amount of internal force needed at the the equilibria. In fact, for typical values of the aerodynamic and kinematic parameters the difference in the amount of internal forces generated in PJD1 design remains small compared to PJD2 case, as it can be seen for example in the simulations of Section V. This supports the viability of the concept from a mechanical perspective by using the simpler design PJD1, if one does not want to resort to the particular joint design in PJD2.
@@ IV Stabilization of the Equilibria
This section presents the solution of the problem of whether every equilibrium point in the set $\mathbb{D}^{[i]}_{x_{e}}$, given in (11), can be made closed-loop asymptotically stable, at least locally, by an action of a regular state feedback law. For clarity, we first provide an informal outline of the solution’s derivation and establish necessary definitions, leading to the core findings stated in Theorem 1 and Proposition 2.
IV-A Informal Outline of the Solution
@@
Denote with $\Sigma_{i}(x;u)$ a system $\Sigma_{i}$ whose states and inputs are respectively $x$ and $u$. The analysis departs by noticing that there exists no static state feedback control $u=\alpha(x)$ rendering the closed-loop system $\Sigma_{1}(x;\alpha(x))$ in (4) locally asymptotically stable (LAS) in an open neighborhood $\mathbb{U}_{{x}_{e}}\subseteq\mathbb{X}$ of an equilibrium ${x}_{e}\in\mathbb{D}^{[i]}_{x_{e}}$, since it turns out that $\Sigma_{1}$ does not possess a well-defined vector relative degree at any $x_{e}\in\mathbb{D}^{[i]}_{x_{e}}$  [22]. However, the goal of stabilizing this closed-loop system at $x_{e}$ is still attainable by means of dynamic state feedback.
Towards this end, we first observe that $\Sigma_{1}$ in (4) can be decomposed into two subsystems
$\Sigma_{1}:\Sigma_{2}(x_{12},x_{2},q_{5};u,x_{11})\text{ and }\Sigma_{3}(x_{11};x_{21})$ (13)
where $\Sigma_{2}$ is found to belong to the class of dynamically input/output feedback linearizable systems w.r.t an output taken from $V_{1}^{0}=:x_{21}$. Second, by applying the dynamic extension algorithm [23] on $\Sigma_{2}$, an extended system $\Sigma_{\bar{2}}(\bar{x};w,x_{11})$ results, where the extended state is $\bar{x}\in\bar{\mathbb{X}}\equiv\mathbb{T}^{8}\times se(3)\times\mathbb{R}^{10}$ and $w\in\mathbb{R}^{6}$ is the new input for its input precompensator. For this extended system $\Sigma_{\bar{2}}$, we compute a local diffeomorphism $H(\bar{x}):\mathbb{U}_{\bar{x}_{e}}\rightarrow\mathbb{U}_{\tilde{x}_{e}}$, with a domain in $\mathbb{U}_{\bar{x}_{e}}\subseteq\bar{\mathbb{X}}$ of $\bar{x}_{e}\in\mathbb{D}^{[i]}_{x_{e}}\cup\mathbb{D}^{[i]}_{f_{e}}$ (12), transforming it into its normal form represented by $\Sigma^{H}_{\tilde{2}}(\zeta,\eta;w,x_{11})$.
This normal form $\Sigma^{H}_{\tilde{2}}(\tilde{x};w,x_{11})$ has a state $\tilde{x}:=(\zeta,\eta)\in\mathbb{U}_{\tilde{x}_{e}}\subseteq\tilde{\mathbb{X}}\equiv se(3)\times se(3)\times\mathbb{T}^{6}\times\mathbb{R}^{6}$, where $\mathbb{U}_{\tilde{x}_{e}}$ denotes an open neighborhood of $\tilde{x}_{e}=:({\zeta_{e}},{\eta_{e}})=H(\bar{x}_{e})$. The normal form is a feedback interconnection between the dynamics of the I/O map $\Sigma_{\zeta}(\zeta;\eta,w,x_{11})$ and some internal dynamics $\Sigma_{\eta}(\eta;\zeta,x_{11})$. $\Sigma_{\zeta}$ can be feedback linearized, becoming $\Sigma_{\zeta}^{w}(\zeta;v)$, by an action of a regular static state feedback $w=\alpha_{w}(\tilde{x},x_{11},v)$ with a domain in $\mathbb{U}_{\tilde{x}_{e}}\times SE(3)\times se(3)$, where $v\in se(3)$ is a virtual input.
At this point, we return to the $\Sigma_{1}$ decomposition in (13). Let us denote with $\Sigma^{H}_{3}(x_{11};\zeta)$ the system $\Sigma_{3}$ in (13) after the diffeomorphism $H$ is applied and with $\Sigma_{\bar{\zeta}}(\bar{\zeta};v)$ its composition with the feedback linearized system $\Sigma_{\zeta}^{w}$, i.e. $\Sigma_{\bar{\zeta}}$ is given by
$\Sigma_{\bar{\zeta}}:\Sigma_{\zeta}^{w}(\zeta;v)\text{ and }\Sigma^{H}_{3}(x_{11};\zeta)$ (14)
where $\bar{\zeta}$ denotes $\zeta$ extended with $x_{11}$. We can now conclude that if there exists a static state feedback $v=\alpha_{v}(\bar{\zeta})$ locally asymptotically stabilizing $\Sigma_{\bar{\zeta}}$ at $\bar{\zeta}_{e}:=(\zeta_{e},x_{11e})$, then, from Theorem 10.3.1 in [22], the cascaded interconnection of the autonomous system $\Sigma_{\bar{\zeta}}(\bar{\zeta};\alpha_{v}(\bar{\zeta}))$ and $\Sigma_{\eta}(\eta;\bar{\zeta})$ is guaranteed to be LAS in $\mathbb{U}_{\tilde{x}_{e}}\times SE(3)$ of $(\tilde{x}_{e},x_{11e})$, provided that the zero dynamics $\Sigma_{\eta}(\eta;0)$ is LAS at $\eta_{e}$}.
This means that, as long as trajectories of $\Sigma_{\bar{\zeta}}$ and $\Sigma_{\eta}$ remain in $\mathbb{U}_{\tilde{x}_{e}}\times SE(3)$, any base pose $x_{11e}$ of the original system $\Sigma_{1}$ (4) can be made locally asymptotically stable by a dynamic state feedback.
IV-B Main Results
Theorem 1 formalizes this result while Proposition 2 provides the stabilizing feedback $v=\alpha_{v}(\bar{\zeta})$ whose construction details, drawing upon a backstepping strategy on $\Sigma_{\bar{\zeta}}$, are laid out in the proof in Sec. VII-D.
@@ IV-B Main Results
Let the original system $\Sigma_{1}$ (4) be decomposed as in (13) and define the system $\Sigma_{3}(x_{11};x_{21})$ to be $\Sigma_{3}:\dot{x}_{11}=[{x}_{21}]\,{x}_{11}$ while $\Sigma_{2}(x_{12},x_{2},q_{5};u,x_{11})$ is
$\Sigma_{2}:\left\{\begin{aligned} \dot{x}=&\left(\begin{array}[]{c}x_{22}\\ M^{-1}(-h-g+\bar{\tau}_{f})\\ 0_{2\times 1}\end{array}\right)+\left(\begin{array}[]{c}0_{6\times 6}\\ M^{-1}\left(\begin{array}[]{cc}D_{b}&0_{6\times 2}\\ D_{t}&0_{6\times 3}\\ \end{array}\right)\\ \begin{array}[]{cc}0_{2\times 4}&I_{2\times 2}\end{array}\end{array}\right)u\\ =&F(x)+G(x)u\end{aligned}\right$.
Then, these statements are true
@@
    $\Sigma_{2}$ is dynamically input/output feedback linearizable w.r.t the I/O pair $(u,h_{y})$, in an open neighborhood $\mathbb{U}_{\bar{x}_{e}}\subseteq\bar{\mathbb{X}}$ of $\bar{x}_{e}\in\mathbb{D}^{[i]}_{x_{e}}\cup\mathbb{D}^{[i]}_{f_{e}}$, where the output function $h_{y}(t):t\rightarrow se(3)$ is $h_{y}(t)=V_{1}^{0}=:x_{21}$. This is accomplished by the input precompensator
    $z(t)=col(z_{i}):=f(t)\in\mathbb{R}^{4},\dot{z}=\bar{w}\in\mathbb{R}^{4},i=2,\ldots,5$ (16)
    and the linearizing state feedback ${w}:=(\bar{w},u_{v_{1}},u_{v_{2}})^{T}$,
    ${w}=D_{2}^{-1}M_{b}(-F_{2}+v)$ (17)
    where $v\in se(3)$ is the virtual input, and $F_{2}$, $D_{2}$ and $M_{b}$ are given in (68), (61) and (66). This path-connected open neighborhood $\mathbb{U}_{\bar{x}_{e}}$ of $\bar{x}_{e}$ is defined as
@@
    For any base pose $x_{11}\in SE(3)$, The original system $\Sigma_{1}$, extended with the precompensator (16), has locally, in $\mathbb{U}_{\bar{x}_{e}}\times SE(3)$, asymptotically stable (LAS) zero dynamics w.r.t the output $h_{y}$, iff $\forall\bar{x}\in\mathbb{Z}_{x}-\{\bar{x}_{e}\}$
    $\|\tilde{f}\|<\frac{1}{\|D_{3}(q_{t}-q_{te})\|}\lambda_{\text{min}}(D_{f})\|\dot{q}_{t}\|$ (19)
    where $\tilde{f}=col({f}_{i})\,,i=2,3,4$, $\mathbb{Z}_{x}$ is a set defined in (77), $D_{3}$ is given in (71) and $D_{f}>0\in\mathbb{R}^{6\times 6}$ contains the constant passive joints friction coefficients.
@@
See Sec. VII-C

### v2

@@ III Study of Attainable Equilibria and
In this section we obtain the set of forced equilibria of the dynamics (1) and investigate their properties in terms of the corresponding four thrust forces produced at these configurations. We begin by giving some preliminary definitions. Afterwards, the main result, which establishes the Equilibria property and the Internal-force property, is summarized in Proposition 1 whose proof is found in Sec. -D of [29].
Imposing $\dot{x}{(t)}=0$ in (1) implies $({\color[rgb]{0,0,0}{V_{1}^{0},\dot{q}_{p}}},{\color[rgb]{0,0,0}{\dot{q}_{a}}})=0$, which defines an equilibrium condition at which the state and the input become, respectively, $x_{e}=(C^{0}_{1e},{\color[rgb]{0,0,0}{q_{pe}}},0,0,{\color[rgb]{0,0,0}{q_{ae}}})^{T}$ and $u_{e}=(f_{e}\,\,u_{ve})^{T}$, with $f_{e}=(f_{2e}\,\,f_{3e}\,\,f_{4e}\,\,f_{5e})^{T}$ and $u_{ve}=(u_{v_{1e}}\,\,u_{v_{2e}})^{T}$. Under this condition, both $h({\color[rgb]{0,0,0}{C_{1}^{0},q_{p}}},{\color[rgb]{0,0,0}{V_{1}^{0},\dot{q}_{p}}})=0$ and $\bar{\tau}_{f}({\color[rgb]{0,0,0}{\dot{q}_{p}}})=0$ vanish, as they depend quadratically and linearly, respectively, on the configuration velocity.
For PJD1, (1) reduces to this static balance vector equation whose admissible solutions are the equilibrium state and input pairs $(x_{e},u_{e})$ such that
$g_{b}(C^{0}_{1e},{\color[rgb]{0,0,0}{q_{pe}}})$ $=(\,\,{\color[rgb]{0,0,0}{D_{b}}}\,\,0_{6\times 2}\,\,)u_{e},$ (3)
${\color[rgb]{0,0,0}{g_{p}}}(C^{0}_{1e},{\color[rgb]{0,0,0}{q_{pe}}})$ $=(\,\,{\color[rgb]{0,0,0}{D_{p}}}\,\,0_{6\times 3}\,\,)u_{e}$
$=(\,\,\tau_{d2}\,\,0\,\,\tau_{d3}\,\,0\,\,\tau_{d4}\,\,0\,\,)_{6\times 1}^{T}$
$0_{2}$ $=u_{ve}$ (4)
where $\tau_{dj}=k_{j}f_{je}\sin(q_{j2e})$, with $j=2,3,4$. $C^{0}_{1e}$ and ${\color[rgb]{0,0,0}{q_{pe}}}=\operatorname{col}(\,q_{je})$, with $q_{je}=\operatorname{col}(\,q_{jie})$ $\forall j=2,3,4,\forall i=1,2$.
@@ III-A Effect of the Aerodynamic Drag Torque at the Equilibrium
The presence of the torque $\tau_{dj}$ in (4) leads to the important consequence that the value of the joints $\color[rgb]{0,0,0}{q_{pe}}$} at the equilibrium is not solely determined by the gravitational torque at equilibrium $\color[rgb]{0,0,0}{g_{pe}}$}, as it is the case for the 2D MAV presented in  [6], but also by a component of the equilibrium thrust $f_{e}$ at equilibrium.
This residual torque $\tau_{dj}$ can be eliminated at any equilibrium by the spring introduced in PJD2 whose proposed implementation in Fig. 2 enables a torsion spring to counteract residual torque and passively return the link to its vertical configuration, zeroing the internal force as stated in proposition 1. In fact, we can tune each variable-stiffness torsion spring such that its stiffness $k_{s}=\frac{\partial\tau_{sj}}{\partial q_{j2}}$ varies linearly with $\cos(q_{j2})$, with a constant slope $\bar{k}=k_{j}f_{je}$, i.e. $k_{s}=\bar{k}\cos(q_{j2})$, which is feasible since $f_{je}$ is found to be constant for any equilibrium of PJD2 as long as the vehicle weight remains constant during flight as shown in Sec. -D of [29]. Hence, the spring needs not to change or be re-tuned between flights. As a result, this component $\tau_{dj}$ is canceled at any PJD2 equilibrium, i.e. when $\tau_{dj}=-\tau_{sje}$, transforming (4) to
${\color[rgb]{0,0,0}{g_{p}}}(C^{0}_{1e},{\color[rgb]{0,0,0}{q_{pe}}})=0_{6},\,\,\,u_{ve}=0_{2}.$ (5)
@@ III-B Minimal Thrust at Equilibrium vs Presence of Internal Forces
Before presenting the result on the Equilibria and Internal-force properties of the proposed MAV (1) we provide two definition for the thrust magnitudes at the equilibrium, in the case in which they are minimal and in the case in which there is internal force.
@@
$\mathcal{F}_{\rm min}^{\rm eq}:=\bigg\{f=col(f_{i})\in\mathbb{R}_{\geq 0}^{l}\bigg|\,\sum^{l}_{i=1}{f}_{i}=m_{\rm tot}\,g_{r}\bigg\}$ (6)
where $m_{\rm tot}$ is the robot’s mass and $g_{r}=9.81m/s^{2}$. Instead, $f$ is said to have internal forces at the equilibrium if it belongs to $\mathcal{F}_{\rm int.for.}^{\rm eq}$, defined as:
$\mathcal{F}_{\rm int.for.}^{\rm eq}:=\bigg\{f=col(f_{i})\in\mathbb{R}_{\geq 0}^{l}\bigg|\,\sum^{l}_{i=1}{f}_{i}>m_{\rm tot}\,g_{r}\bigg\}$ (7)
@@
In other words, a minimal thrust magnitudes vector $f_{e}\in\mathcal{F}_{\rm min}^{\rm eq}$ has the least amount of total thrust magnitudes necessary to counteract the gravity at the equilibrium, whereas $f_{e}\in\mathcal{F}_{\rm int.for.}^{\rm eq}$ means that the vehicle generates more total thrust than what is strictly necessary for compensating gravity, thus having some wasted internal force. Moreover, there exists a unique orientation for each propeller in an MAV such that when all the propellers assume such orientation the absence of internal force is guaranteed at hovering, proven in Lemma 1 in Sec. -C [29]. The following proposition establishes the main result of this section.
@@
    For any base pose $C^{0}_{1e}\in SE(3)$, there exists an equilibrium for PJD2 in which the joint equilibrium configuration ${\color[rgb]{0,0,0}{q_{pe}}}^{[2]}$ satisfies $C^{0}_{j}(C^{0}_{1e},q^{[2]}_{je})e_{3}=e_{3}$ $\forall j\in\{2,\dots,5\}$ and a thrust magnitudes vector $f^{[2]}_{e}\in\mathcal{F}_{\rm min}^{\rm eq}$ is produced whose elements $f^{[2]}_{je}={g_{r}(\frac{1}{4}m_{b}+m_{j})}$ are constants $\forall C^{0}_{1e}$ with $k_{i}$ chosen by $\sum_{j=2}^{5}k_{i}{(\frac{1}{4}m_{b}+m_{j})}=0$.
@@
    For any base pose $C^{0}_{1e}$, PJD2 has no internal force at any equilibrium ${\color[rgb]{0,0,0}{q_{pe}}}^{[2]}$, by Definition 1, as a corollary of $f^{[2]}_{e}\in\mathcal{F}_{\rm min}^{\rm eq}$ (Statement 1).
@@
    When $\tau_{dj}\neq 0$, PJD1 has no joint equilibrium ${\color[rgb]{0,0,0}{q_{pe}}}^{[1]}$ in which each propeller produces a minimal thrust: $C^{0}_{j}(C^{0}_{1e},q^{[1]}_{je})e_{3}\neq e_{3},\,\forall\tau_{dj}\neq 0$.
@@
    When $\tau_{dj}\neq 0$, it follows from Statement 3, and Lemma 1 in [29], that PJD1 has an internal force $f^{[1]}_{e}\in\mathcal{F}_{\rm int.for.}^{\rm eq}$. The amount of such force can be reduced by changing parameters $k_{i}$,c\,c and $m_{j}$.
@@
See Sec. -D of [29]. ∎
The state equilibria set for PJDi is denoted by $\mathbb{D}^{[i]}_{x_{e}}$ whereas $\mathbb{D}^{[i]}_{f_{e}}$ signifies their associated thrust equilibrium sets. The explicit relations are given in Sec. -D4 [29].
We conclude the section remarking that the perfect vertical alignment of propellers in any PJD2 equilibrium is possible thanks to each spring torque $\tau_{sje}$ absorbing the residual drag $\tau_{dj}$ affecting the joint variable $q_{j1}$. This allows any joint equilibrium ${\color[rgb]{0,0,0}{q_{pe}}}^{[2]}$ to be dictated only by the gravity acting on the CoMs of the pendulum-like links B2-B4, given that the pivots, i.e., the joints, are passive.
@@
The design using PJD1 constitutes still an excellent compromise between mechanical simplicity and energy efficiency, with the ability to outperform the state of the art designs when it comes to the amount of internal force needed at the the equilibria. In fact, for typical values of the aerodynamic and kinematic parameters, the difference in the amount of internal forces generated in PJD1 design remains significantly small compared to the ideal PJD2 case, as it can be seen for example in the simulations of Section V. This supports the viability of the concept from a mechanical perspective by using the simpler design PJD1, if one does not want to resort to the particular joint design in PJD2.
@@ IV Stabilization of the Equilibria
Following the characterization of the Equilibria and Internal-Force properties in Proposition 1, this section presents the solution of the problem of whether every equilibrium point in the equilibria set $\mathbb{D}^{[i]}_{x_{e}}$ can be made closed-loop asymptotically stable, at least locally, by a regular state feedback law, hence establishing the Stabilizability property. The core findings are stated in Theorem 1 and Proposition 2. The proposed control architecture is shown in Fig. 3.
@@
The analysis departs by noticing that there exists no static state feedback control law rendering the closed-loop system $\Sigma_{1}$ in (1) locally asymptotically stable (LAS) in an open neighborhood $\mathbb{U}_{{x}_{e}}\subseteq\mathbb{X}$ of an equilibrium ${x}_{e}\in\mathbb{D}^{[i]}_{x_{e}}$, since it turns out that $\Sigma_{1}$ does not possess a well-defined vector relative degree at any $x_{e}\in\mathbb{D}^{[i]}_{x_{e}}$  [23]. However, the goal of stabilizing this closed-loop system at $x_{e}$ remains attainable by a dynamic state feedback obtained from the dynamic extension algorithm [24], as illustrated in Theorem 1.
Figure 3: Block diagram of the proposed control scheme. It takes a desired base pose $C^{0}_{1e}\in SE(3)$ and converts it into a reference base twist $\zeta_{1e}(t)$ through configuration error feedback $E_{c}$. Dynamic FBL is applied to the twist dynamics, which are stabilized using a PID controller. The accelerations are efficiently computed via a modified Articulated-Body Inertia Algorithm (ABA) in Sec. -B of [29].
@@
Let the original system $\Sigma_{1}$ (1) be decomposed into $\Sigma_{2}$ and $\Sigma_{3}$, respectively given by $\Sigma_{3}:\dot{C}_{1}^{0}=[V_{1}^{0}]\,C_{1}^{0}$ and this $\Sigma_{2}$
$\Sigma_{2}:\left\{\begin{aligned} \dot{x}_{2}=&{\color[rgb]{0,0,0}{F_{2}(x)+G_{2}(x)u}}\\ =&\left(\begin{array}[]{c}{\color[rgb]{0,0,0}{\dot{q}_{p}}}\\ M^{-1}(-h-g+\bar{\tau}_{f})\\ 0_{2\times 1}\end{array}\right)+\left(\begin{array}[]{c}0_{6\times 6}\\ M^{-1}\left(\begin{array}[]{cc}D_{b}&0_{6\times 2}\\ D_{t}&0_{6\times 3}\\ \end{array}\right)\\ \begin{array}[]{cc}0_{2\times 4}&I_{2\times 2}\end{array}\end{array}\right)u\end{aligned}\right$.
Then, these statements are true:
@@
    $\Sigma_{2}$ is dynamically input/output feedback linearizable w.r.t the I/O pair $(u,h_{y})$, where the output function $h_{y}(t):t\rightarrow\mathbb{R}^{6}$ is $h_{y}(t)={\color[rgb]{0,0,0}{V_{1}^{0}}}$, by a dynamic extension [24] given by this input precompensator
    $z(t)=col(z_{i}):=f(t)\in\mathbb{R}^{4},\dot{z}=\bar{w}\in\mathbb{R}^{4},i=2,\ldots,5.$
    The extended state is $\bar{x}=({\color[rgb]{0,0,0}{q_{p}}},{\color[rgb]{0,0,0}{V_{1}^{0},\dot{q}_{p}}},{\color[rgb]{0,0,0}{q_{a}}},z)\in\bar{\mathbb{X}}\equiv\mathbb{T}^{8}\times\mathbb{R}^{16}$ whereas the linearizing state feedback on the extended system, whose I/O pair is $(w,h_{y})$, is given by
    ${w}=\left(\begin{smallmatrix}\bar{w}\\ u_{v_{1}}\\ u_{v_{2}}\end{smallmatrix}\right)=D_{2}^{-1}M_{b}(-E_{2}+v)$ (10)
    where $v\in\mathbb{R}^{6}$ is the virtual input, and $E_{2}$, $D_{2}$ and $M_{b}$ are given in (62), (55) and (60) of Sec. -E [29]. This dynamic FBL is well-defined in an open neighborhood $\mathbb{U}_{\bar{x}_{e}}\subseteq\bar{\mathbb{X}}$ of $\bar{x}_{e}\in\mathbb{D}^{[i]}_{x_{e}}\cup\mathbb{D}^{[i]}_{f_{e}}$ defined as
@@
    For any base pose ${\color[rgb]{0,0,0}{C_{1e}^{0}}}\in SE(3)$, The original system $\Sigma_{1}$, extended with the precompensator (9), has locally, in $\mathbb{U}_{\bar{x}_{e}}\times SE(3)$, asymptotically stable (LAS) zero dynamics w.r.t the output $h_{y}$, iff $\forall\bar{x}\in\mathbb{Z}_{x}-\{\bar{x}_{e}\}$
    $\|\tilde{f}\|<\frac{1}{\|D_{3}({\color[rgb]{0,0,0}{q_{p}}}-{\color[rgb]{0,0,0}{q_{pe}}})\|}\lambda_{\text{min}}(D_{f})\|{\color[rgb]{0,0,0}{\dot{q}_{p}}}\|$ (12)
    where $\tilde{f}=col({f}_{i})\,,i=2,3,4$, $\mathbb{Z}_{x}$ and $D_{3}$ are respectively given in (71) and (65) in [29]. $D_{f}>0\in\mathbb{R}^{6\times 6}$ contains the constant passive joints friction coefficients.
@@
See Sec. -E in [29].∎
This means that, as long as the extended system trajectories remain in $\mathbb{U}_{\bar{x}_{e}}\times SE(3)$ where the decoupling holds, any base pose $C^{0}_{1e}$ of the original system $\Sigma_{1}$ (1) can be made locally asymptotically stable by a dynamic state feedback. Proposition 2 constructs the the virtual controller that stabilizes the resulting linearized twist dynamics.